Dianne was given $35,000 when she retired. she invested some at 7% interest and the rest at 9% interest. if she earned $2,910 in interest in one year, how much was invested in each account?
Answer by checkley77(12844)

.09x+.07(35,000-x)=2,910
.09x+2,450-.07x=2,910
.02x=2,910-2,450
.02x=460
x=460/.02
23,000 invested @ 8%
35,000-23,000=12,000 invested @ 7%
Proof:
.09*23,000+.07*12,000=2,910
2,070+840=2,910
1,910=2,910